OpenStudy (anonymous):

I don't understand Geometric series and this problem. Evaluate S_5 for 600+300+150+.... Choices 1,162.5 581.25 18600 37.5 I got the wrong answer twice and it's not 37.5 Please help...

jhonyy9 (jhonyy9):

so than you see first term is 1. 600 and 2. 600/2 = 300 than 3.300/2=150 4.150/2=75 5. 75/2=37,5 this is clearly now ?

OpenStudy (dayakar):

a1=600 a2= 300 a3 = 150 common ratio (r) = a3/a2 = 150/300 = 1/2 [ r is less than 1] sn = a {r^n -1]/(r-1) [ here a= a1= 600, r= 1/2 , n= 5] find s5=